OpenStudy (kittykat):

Compute u × v where u = (1, 2, 3) and v = (1, 1, 2).

OpenStudy (daniyil_the_spy):

ok

OpenStudy (caozeyuan):

2*2-1*3)i-(1*2-1*3)j+1*1-1*2)k

OpenStudy (caozeyuan):

so 1i+1j-1k

OpenStudy (kittykat):

wait is there are formula when did you do 2*2 first

OpenStudy (mww):

you need to understand determinants first which always pops up when you do cross products.

OpenStudy (kittykat):

yes i know that but what are the steps u are using

OpenStudy (mww):

http://www.mathsisfun.com/algebra/matrix-determinant.html It's tricky to explain directly but you pick the i element, then cross out then row and column with it (think Sudoku elimination) then find the determinant of the remaining 2x2 matrix, do the same for element j and k. Signs interchange however so you have i det1 - j det 2 + k det 3.

OpenStudy (kittykat):

I dont get how the middle one is plus 1 it should be negative 1 right ? since you did 1*2-1*3

OpenStudy (kittykat):

which would be 2-3=-1

OpenStudy (kittykat):

oh wait cuz its the rule always do + then - then + right?

OpenStudy (mww):

yeah + - + etc. it alternates

OpenStudy (kittykat):

Find a vector equation of the line passing through P(3, −1, 4) and perpendicular to the plane 3x − 2y − z = 0.

OpenStudy (kittykat):

can u help me with this one

OpenStudy (mww):

A normal vector to the plane of the form ax + by + cz = d is (a, b, c) So for 3x - 2y - z = 0 a normal vector is (3,-2,-1) You just need your line to be parallel to this normal vector and passing through the point P In general we write Q(v) = P(v) + A(v) t where P(v) is the vector pt Q passes through and A(v) is your parallel vector to Q(v) and t is your parameter.

OpenStudy (kittykat):

Find the equations of the line of intersection of the planes 3x + y − 2z = 1 and x + y + z = 5.

OpenStudy (kittykat):

can you help me with one too

OpenStudy (caozeyuan):

a line is a point and a direction vector

OpenStudy (caozeyuan):

since the point is on both plane 1 and plane 2, you can solve the system of equations for a particular solution

OpenStudy (caozeyuan):

(0, 7/3,4/3) is a good one, you can verify that this does satisfy both of the euqations

OpenStudy (caozeyuan):

y is 11/3, sorry

OpenStudy (caozeyuan):

the direction vector is perpendicular to the normal of plane 1 and 2

OpenStudy (caozeyuan):

therefore it is the cross of the two normals

OpenStudy (caozeyuan):

(3,1,-2)x(1,1,1)=(3,-5,2)

OpenStudy (caozeyuan):

answer is (0,11/3,4/3)+(3,-5,2)*t